# Meta-learning-based percussion transcription and $t\bar{a}la$ identification from low-resource audio

### v2

@@ Abstract
This study introduces a meta-learning-based approach for low-resource Tabla Stroke Transcription (TST) and $t\bar{a}la$ identification in Hindustani classical music. Using Model-Agnostic Meta-Learning (MAML), we address the challenge of limited annotated datasets, enabling rapid adaptation to new tasks with minimal data. The method is validated across various datasets, including tabla solo and concert recordings, demonstrating robustness in polyphonic audio scenarios. We propose two novel $t\bar{a}la$ identification techniques based on stroke sequences and rhythmic patterns. Additionally, the approach proves effective for Automatic Drum Transcription (ADT), showcasing its flexibility for Indian and Western percussion music. Experimental results show that the proposed method outperforms existing techniques in low-resource settings, significantly contributing to music transcription and studying musical traditions through computational tools.
@@ I Introduction
Music Information Retrieval (MIR) focuses on extracting meaningful features from music recordings, enabling diverse applications in music education, production, and musicology [1]. Among its many tasks, Automatic Music Transcription (AMT) is particularly challenging, as it involves converting audio recordings into symbolic representations such as scores or MIDI files. AMT supports critical tasks such as melody extraction, chord recognition, beat tracking, and rhythm analysis, forming the foundation for advanced music research and applications [2]. While substantial progress has been made in transcribing melodic instruments like the piano and guitar, the transcription of percussive instruments remains under-explored [1]. Furthermore, research progress in the percussion transcription of Hindustani classical music has been significantly slower compared to Western music [3].
@@
The tabla is a key instrument in Hindustani music, serving as the rhythmic backbone and accompanying singers and instrumentalists. It helps maintain the structure and essence of the chosen raga (melodic framework). Tabla Stroke Transcription (TST) is crucial for analysing the complex rhythmic patterns ($t\bar{a}las$) that form the basis of Hindustani music. Accurate TST also aids in preserving and digitising traditional music and creating learning tools for learners and musicians. Like Automatic Drum Transcription (ADT) in Western music, which focuses on drum sound transcription to understand rhythm structures, organise music collections and develop advanced tools, TST aims to identify and classify rhythmic events. However, unlike ADT, TST presents distinct challenges due to the tabla’s unique physical and timbral characteristics compared to drums [1, 4].
@@
TST and ADT are challenging due to the complexity of audio signals in music tracks, where multiple instruments (polyphony) often mask rhythmic sounds. In Western music, bass guitars may overlap with bass drums, while melodic instruments or vocals can obscure tabla strokes in Hindustani music. Transcription methods address these challenges at different levels of complexity: Tabla/Drum-only Transcription (DTD) focuses on isolated recordings, avoiding polyphony. Tabla/Drum Transcription with Percussion (DTP) includes additional percussive elements, adding complexity. The most challenging, Tabla/Drum Transcription with Melodic Instruments (DTM), handles full audio mixes with all instruments and vocals. Since DTM represents most real-world musical tracks, it is the most critical and complex task in transcription.
@@
Most of the present works in ADT rely on deep learning (DL) models due to their ability to handle complex audio mixtures. Similarly, few studies have been conducted on TST, with recent efforts also focusing on DL models [5, 6]. However, DL models require large, annotated datasets for supervised training, which are scarce. Creating these datasets is challenging because annotations must be precisely aligned with the audio, a labour-intensive process requiring skilled musicians. Additionally, many annotated tracks are copyrighted, limiting their public availability. To address this, synthetic data generated from MIDI tracks has been used to produce large-scale training datasets with accurate labels. Despite their large volume, synthetic datasets often perform less effectively than real-world annotated data, making dataset scarcity a significant challenge for DL-based ADT and TST. To overcome this limitation, this paper introduces a new stroke transcription method based on model-agnostic meta-learning (MAML) [7], which operates effectively with minimal data.
TST provides tabla strokes, and $t\bar{a}la$ is typically identified by stroke sequence. While some studies have explored music sequence matching [8, 9], these approaches are not directly applicable to $t\bar{a}la$ identification. In this study, we propose two new $t\bar{a}la$ identification methods.
The key contributions of this work are summarised as follows:
@@
    A novel meta-learning-based approach for Tabla Stroke Transcription (TST) is proposed and validated on both tabla solo and complete concert recordings. This method is further validated for Automatic Drum Transcription (ADT) tasks using DTD, DTP, and DTM datasets.
@@
The rest of the paper is organized as follows: Section II provides a brief overview of the related literature. Section III introduces the fundamental concepts of $t\bar{a}la$ in Hindustani music. Section IV provides a detailed description of the proposed method for TST, followed by $t\bar{a}la$ identification, and finally, ADT. Section V outlines the experimental setup for evaluating the proposed approach, including the datasets used for TST, $t\bar{a}la$ identification, and ADT. Section VI discusses the experimental evaluation and findings, and Section VII concludes the paper with key insights and future directions.
@@ II-A TST and ADT
Significant work has been done on ADT in Western music [1]. Recent ADT methods utilize deep learning with various model architectures and feature representations [10, 11, 12, 1, 13, 14, 15, 16, 17, 18, 19, 2]. ADT identifies stroke onsets, marked by sudden energy increases. The attack-decay-sustain-release (ADSR) sequence forms the general temporal structure of a stroke. Detecting strokes in a drum kit is relatively straightforward due to distinct timbral structures and variations in shape and material [1]. Drum strokes feature sharp attacks, fast decay, and uniform repetition [20]. In contrast, TST is more challenging than ADT because of the significant timbral similarities and irregular shapes of strokes produced by the same instruments. This complexity requires sophisticated techniques for accurate transcription, making rhythmic analysis in Hindustani music more intricate. This work focuses on TST, emphasizing $t\bar{a}la$ identification and further validation of TST methods alongside ADT.
The literature on TST employs two main approaches: segment-classify and deep learning. The segment-classify method detects stroke onsets using thresholding on spectral flux, then classifies segments with various classifiers. Studies [21, 22, 23, 24, 25, 26, 27] follow this approach. Early work used Gaussian Mixture Models (GMM) and Hidden Markov Models (HMM), later extended with spectral and temporal features for neural networks and SVM classifiers [21]-[23]. Studies [24]-[27] focused on acoustic features and ADSR properties with MLP and SVMs. The DL approach involves spectrograms or Mel spectrograms processed through networks for transcription, with recent studies [9, 5, 6] employing CNNs, data augmentation, and transfer learning techniques.
Early studies [21, 22, 23, 24, 25, 26, 27] trained models on limited datasets, often from a single tabla. Recent studies [5, 6] used more realistic datasets with harmonium accompaniment but still employed only four stroke classes, insufficient for actual $t\bar{a}la$ identification. Additionally, these models have not been tested on concert data with vocals and other instruments, raising doubts about their performance on actual stroke classes and concert datasets.
@@ II-B Meta-learning
Meta-learning has been extensively explored in image processing and computer vision to enable deep-learning models with minimal data samples. However, its application in the audio domain remains comparatively limited. Meta-learning mimics the human ability to learn from previous experiences or knowledge. Meta-learning algorithms are further divided into metric-based, model-based, and optimization-based approaches. References [28, 29, 30, 31, 32, 33, 34, 35, 36] primarily focus on metric-based few-shot learning, especially using prototypical networks. For instance, [30, 36] applies prototypical networks to drum transcription, treating each instrument as a binary classification task. Similarly, [31, 32, 33] handle individual audio event clips as binary classification tasks. In [34], multi-class audio segments are framed as binary classification, with the target class as positive and others as negative. Furthermore, [35] adapts the MAML algorithm for sound localization in new room environments, while [37] applies it to melody extraction.
@@ II-C $T\bar{a}la$ identification (Music sequence matching)
String-matching techniques have been extensively explored in the literature for comparing two musical fragments. In [8], Longest Common Subsequence (LCS) and Rough Longest Common Subsequence (RLCS) matching methods are presented. In [38], a modified version of RLCS is discussed for detecting motifs within $alapanas$ in Carnatic music. Similarly, [9] and [39] employ a modified RLCS approach for searching specific stroke patterns in tabla stroke transcription. In [40], $R\bar{a}ga$ verification utilises the Longest Common Segment Set (LCSS). However, LCS, RLCS, and LCSS operate on the subsequence principle, where elements occur in the same order but not necessarily contiguously. In the context of $t\bar{a}las$, where continuity and order are crucial, these measures may not be suitable for analysing continuous sequences. References [3] and [41] focus on matching beat amplitude patterns, where a specific stroke corresponds to a particular amplitude. However, this approach may fail in practical situations where beat amplitudes do not consistently follow the same pattern.
Figure 1: System overview for automatic $t\bar{a}la$ identification
III $T\bar{a}la$ in Hindustani Music
In Hindustani music, a $t\bar{a}la$ describes the rhythmic framework of any composition. A $t\bar{a}la$ consists of a fixed time cycle divided into equal primary time units known as $m\bar{a}tra$. These $m\bar{a}tra$ are grouped into sections, sometimes with varying durations, referred to as $vibh\bar{a}gs$. The complete cycle is called an $avart$, and the start of the cycle (the downbeat) is called $s\bar{a}m$. Each $t\bar{a}la$ features pre-defined rhythmic patterns called the $thek\bar{a}$ [42]. Comprehensive explorations of $t\bar{a}la$ in Hindustani music can be found in [43, 42]. The tabla is the main percussion instrument that acts as the timekeeper of $t\bar{a}la$. It comprises two drums: the left-hand bass drum, known as the $bayan$ or $dagga$, and the right-hand drum, referred to as the $dayan$ or $tabla$. These drums can produce various pitched sounds and are played using palms and fingers.
[…]
Even though Hindustani and Western music differ, comparing some terms is useful. In Hindustani music, $m\bar{a}tra$ and $avart$ are similar to ‘beat’ and ‘bar’ in Western music. To clarify things, we’ll use ‘beat’ instead of ‘$m\bar{a}tra$’ and ‘stroke’ instead of ‘$b\bar{o}l$.’ This helps us understand the rhythms better across different musical styles.
IV Proposed Method
This article introduces a MAML-based approach for Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT) with limited data. By leveraging the stroke sequence obtained from TST, we perform $t\bar{a}la$ identification using two proposed methods. The proposed tala identification system is illustrated in Fig. 1, showcasing the integration of these components.
IV-A Tabla Stroke Transcription (TST)
@@
This section presents our solution for tabla stroke transcription with minimal labeled data. We train and test deep learning models using three methods to identify the optimal transcription approach. Subsequently, we leverage the pre-trained model weights from the best method for transfer learning and model-agnostic meta-learning (MAML) for tabla stroke transcription.
IV-A1 Step1: Pre-training
@@
We adopt the method introduced by Rohit et al. [5, 6], expanding the stroke classes from four to ten. Ten distinct CNN models are employed to classify the ten stroke classes. Let $(X_{i},Y_{i})$ represent a single sample pair from a labeled dataset comprising $N$ samples of individual strokes. Here, the input $X_{i}\in\mathbb{R}^{3\times F\times T}$ to the model consists of a 3-channel chunk of 15-time frames $(T)$ extracted from a stack of 3 log-Mel spectrograms. The model $f_{\theta}$, parameterized by $\theta$, predicts the presence of onset for the center frame by leveraging contextual information from surrounding frames. It compares its output against a binary target $Y_{i}\in{0,1}$, where 1 denotes stroke onset, and 0 denotes no onset.
@@
This work proposes a new mridangam and tabla stroke transcription method inspired by Sound Event Detection (SED) tasks [44]. The proposed method utilizes a CRNN for stroke transcription. The labeled dataset of $N$ samples is denoted as $D(X_{i},Y_{i})$, where $i=0,1,…,N$. Here, $X_{i}\in\mathbb{R}^{F\times T}$ is the input feature matrix (standardized log Mel-spectrogram), containing $F$ acoustic features extracted from each frame in each audio recording with $T$ time frames. The labels are one-hot encoded as the target output matrix $Y_{i}\in\left\{0,1\right\}^{C\times T}$, where $C$ represents the number of predefined stroke classes.
@@
Post-processing is applied to these predictions in two steps: First, if the predicted class label of $frame_{i-1}$ and $frame_{i+1}$ are the same but different from the class label of frame $frame_{i}$ then the class label of $frame_{i}$ is changed to match $frame_{i-1}$ and $frame_{i+1}$. In the second step, the frame where the class label changes from one class to another is identified as the onset frame for the new class. All onset frames in the test sample are identified, and the temporal locations of the onsets are calculated from these frames.
@@
[…]
IV-A2 Step 2 - A): Model Agnostic Meta-learning (MAML)

### v3

@@ Abstract
This study introduces a meta-learning-based approach for low-resource Tabla Stroke Transcription (TST) and $t\bar{a}la$ identification in Hindustani classical music. Using Model-Agnostic Meta-Learning (MAML), we address the challenges of limited annotated datasets and label heterogeneity, enabling rapid adaptation to new tasks with minimal data.
The method is validated across various datasets, including tabla solo and concert recordings, demonstrating robustness in polyphonic audio scenarios. We propose two novel $t\bar{a}la$ identification techniques based on stroke sequences and rhythmic patterns. Additionally, the approach proves effective for Automatic Drum Transcription (ADT), showcasing its flexibility for Indian and Western percussion music. Experimental results show that the proposed method outperforms existing techniques in low-resource settings, significantly contributing to music transcription and studying musical traditions through computational tools.
@@ I Introduction
Music Information Retrieval (MIR) focuses on extracting meaningful features from music recordings, enabling diverse applications in music education, production, and musicology [1]. Among its many tasks, Automatic Music Transcription (AMT) is particularly challenging, as it involves converting audio recordings into symbolic representations such as scores or MIDI files. AMT supports critical tasks such as melody extraction, chord recognition, beat tracking, and rhythm analysis, forming the foundation for advanced music research and applications [2]. While substantial progress has been made in transcribing melodic instruments such as the piano, guitar, strings, and winds, the transcription of percussive instruments remains relatively under-explored [1, 3]. Furthermore, research progress in the percussion transcription of Hindustani classical music has been significantly slower compared to Western music [4].
@@
The tabla is a key instrument in Hindustani music, serving as the rhythmic backbone and accompanying singers and instrumentalists. It helps maintain the structure and essence of the chosen raga (melodic framework). Tabla Stroke Transcription (TST) is crucial for analyzing the complex rhythmic patterns ($t\bar{a}las$) that form the basis of Hindustani music. Accurate TST also aids in preserving and digitizing traditional music and creating learning tools for learners and musicians. Like Automatic Drum Transcription (ADT) in Western music, which focuses on drum sound transcription to understand rhythm structures, organize music collections and develop advanced tools, TST aims to identify and classify rhythmic events. However, unlike ADT, TST presents distinct challenges due to the tabla’s unique physical and timbral characteristics compared to drums [1, 5].
@@
TST and ADT are challenging due to the complexity of audio signals in music tracks, where multiple instruments (polyphony) often mask rhythmic sounds. In Western music, bass guitars may overlap with bass drums, while melodic instruments or vocals can obscure tabla strokes in Hindustani music. Transcription methods address these challenges at different levels of complexity: Drum/Tabla -only Transcription (DTD) focuses on isolated recordings, avoiding polyphony. Drum/Tabla Transcription with Percussion (DTP) includes additional percussive elements, adding complexity. The most challenging, Drum/Tabla Transcription with Melodic Instruments (DTM), handles full audio mixes with all instruments and vocals. Since DTM represents most real-world musical tracks, it is the most critical and complex task in transcription.
@@
Most existing ADT and recent TST methods rely on deep learning (DL) models due to their ability to handle complex audio mixtures [6, 7]. However, DL models require large, annotated datasets for supervised training, which are scarce. Creating these datasets is challenging because annotations must be precisely aligned with the audio, a labour-intensive process requiring skilled musicians, and many annotated tracks are copyrighted, limiting their public availability. Synthetic data generated from MIDI tracks has been used to produce large-scale training datasets with accurate labels, but such datasets often perform less effectively than real-world annotated data, making dataset scarcity a significant barrier for DL-based ADT and TST.
Beyond low-resource conditions, a second challenge is label heterogeneity: stroke labels in tabla and drum datasets differ due to performance styles, labelling schemes, and musical traditions. For example, tabla datasets may follow $ghar\bar{a}n\bar{a}$ conventions, while Western drum datasets use distinct ontologies. These divergences lead to task heterogeneity with mismatched label spaces, reducing the reliability of standard transfer learning approaches that assume shared label ontologies and often requiring time-consuming label harmonization. Model-Agnostic Meta-Learning (MAML) enables fast adaptation across tasks with distinct label spaces [8, 9, 10, 11]. This paper presents an MAML-based stroke transcription method addressing data scarcity and label inconsistency.
In this work, we address the challenge of performing accurate Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT) with limited annotated data. Based on the transcribed stroke sequences from TST, we propose methods for $t\bar{a}la$ identification, as $t\bar{a}la$ forms the rhythmic backbone of Hindustani music. Identifying the $t\bar{a}la$ enables a deeper analysis of Hindustani music, and existing music sequence matching approaches [12, 13] are not directly applicable in this context.
The key contributions of this work are summarized as follows:
@@
    A novel meta-learning-based approach for TST is proposed and validated on both tabla solo and full concert recordings. This method is also extended to ADT tasks using the DTD, DTP, and DTM datasets, demonstrating its generalization across datasets.
@@
The rest of the paper is organized as follows: Section II provides an overview of related literature, Section III details the proposed method for TST and ADT, Section IV presents information about the datasets used, Section V outlines the experimental setup, Section VI discusses the experimental evaluation and findings, and Section VII concludes the paper with key insights and future directions. The dataset and supplementary materials are available at https://github.com/madhavlab/2025_drumtrans_rkodag
@@ II-A TST and ADT
Significant work has been done on ADT in Western music [1]. Recent ADT methods utilize deep learning with various model architectures and feature representations [14, 15, 1, 16, 17, 18, 19, 2]. ADT identifies stroke onsets, marked by sudden energy increases. The attack-decay-sustain-release (ADSR) sequence forms the general temporal structure of a stroke. Detecting strokes in a drum kit is relatively straightforward due to distinct timbral structures and variations in shape and material [1]. Drum strokes exhibit sharp attacks, rapid decay, and consistent repetition [20]. In contrast, TST is more complex due to timbral similarities and variable stroke shapes from the same instrument, necessitating more sophisticated transcription methods and making rhythmic analysis in Hindustani music more intricate. This work focuses on TST, emphasizing $t\bar{a}la$ identification and further validation of TST methods alongside ADT.
The literature on TST employs two main approaches: segment-classify and deep learning. The segment-classify method detects stroke onsets using thresholding on spectral flux, then classifies segments with various classifiers. Studies [21, 22, 23, 24] follow this approach. Early work used Gaussian Mixture Models (GMM) and Hidden Markov Models (HMM), later extended with spectral and temporal features for neural networks and SVM classifiers [21, 22, 23]. Study [24] focuses on the acoustic features and ADSR properties of tabla strokes. The DL approach involves spectrograms or Mel spectrograms processed through networks for transcription, with recent studies [25, 6, 7] employing CNNs, data augmentation, and transfer learning techniques.
Early studies [21, 22, 23, 24] trained models on limited datasets, often from a single tabla. Recent studies [6, 7] used more realistic datasets with harmonium accompaniment but still employed only four stroke classes, insufficient for actual $t\bar{a}la$ identification. Additionally, these models have not been tested on concert data with vocals and other instruments, raising doubts about their performance on actual stroke classes and concert datasets.
@@ II-B Meta-learning
Meta-learning, or “learning to learn,” is a training paradigm in which a model is optimized to rapidly adapt to new tasks using limited data [8]. It mimics the human ability to transfer prior experience to new situations by learning how to learn from diverse tasks. Unlike traditional models that train on a fixed dataset and generalize within that domain, meta-learning involves exposure to a distribution of tasks during training, enabling the model to extract transferable knowledge. This knowledge allows the model to quickly fine-tune to new tasks with limited data. The approach typically involves two components: a base-learner that learns task-specific patterns, and a meta-learner that captures cross-task learning strategies to guide fast adaptation [8, 9, 10, 11].
Meta-learning has been extensively explored in image processing and computer vision to enable deep-learning models with limited data samples. However, its application in the audio domain remains comparatively limited. Meta-learning mimics the human ability to learn from previous experiences or knowledge. Meta-learning algorithms are further divided into metric-based, model-based, and optimization-based approaches. References [26, 27, 28, 29, 30, 31, 32] primarily focus on metric-based few-shot learning, especially using prototypical networks. For instance, [27, 32] applies prototypical networks to drum transcription, treating each instrument as a binary classification task. Similarly, [28, 29, 30] handle individual audio event clips as binary classification tasks. In [31], multi-class audio segments are framed as binary classification, with the target class as positive and others as negative. Furthermore, [33] applies the MAML algorithm to melody extraction.
@@ II-C $T\bar{a}la$ Identification (Music Sequence Matching)
In Hindustani music, a $t\bar{a}la$ defines the rhythmic structure of a composition. It comprises a fixed time cycle divided into equal time units called $m\bar{a}tras$, grouped into sections ($vibh\bar{a}gs$), forming a complete cycle known as an $avart$. The starting point of this cycle is referred to as $s\bar{a}m$, and each $t\bar{a}la$ features a characteristic rhythmic pattern known as the $thek\bar{a}$. Comprehensive explorations of $t\bar{a}la$ in Hindustani music can be found in [34, 35]. Despite fundamental differences between Hindustani and Western music, the term equivalence aids clarity. Accordingly, we adopt ‘beat’ for $m\bar{a}tra$ and ‘stroke’ for $b\bar{o}l$ to facilitate cross-cultural rhythmic analysis.
The tabla, consisting of two drums: the left-hand $bayan$ and the right-hand $dayan$, is the primary percussion instrument used to render the $thek\bar{a}$. A wide range of $t\bar{a}las$ exists in Hindustani music, with $T\bar{i}nt\bar{a}l$, $Ek\bar{a}l$, $Jhapt\bar{a}l$, and $Rupak\ t\bar{a}l$ being the most popular and the focus of this study. Six major $ghar\bar{a}n\bar{a}s$: Delhi, Ajrada, Lucknow, Banaras, Farrukhabad, and Punjab, define distinctive tabla styles. These styles vary in stroke articulation, dynamics, ornamentation, and improvisational phrasing, though the structure and number of strokes in the $thek\bar{a}$ typically remain consistent.
Accurate $t\bar{a}la$ identification is essential for tasks such as rhythm analysis, segmentation, and transcription. String-matching techniques have been extensively explored in the literature for music sequence comparison. In [12], Longest Common Subsequence (LCS) and Rough LCS (RLCS) are used for symbolic matching. A modified RLCS approach is presented in [36] for motif detection in Carnatic $al\bar{a}panas$. In the context of tabla, [25]employ a modified RLCS approach for matching specific stroke sequences in transcribed performances. In [37], $R\bar{a}ga$ verification is performed using Longest Common Segment Set (LCSS). While these techniques preserve element order, they do not enforce continuity, limiting their effectiveness for $t\bar{a}la$ patterns where ordered and contiguous strokes are crucial. Alternatively, [4] propose beat amplitude pattern matching, assuming a correlation between strokes and amplitude. However, performance variability in dynamics often renders these methods unreliable in real-world scenarios.
Consequently, there remains a need for matching strategies that retain both continuity and order in stroke sequences while accommodating expressive variations in timing and dynamics.
Figure 1: System overview: (A) TST: Pre-trained on mridangam data using OTM, PTM1, PTM2, and adapted via MAML. (B) $T\bar{a}la$ identification: Identifies $t\bar{a}la$ using transcribed strokes from TST and NW Matching/Stroke Ratio scores. (C) ADT: Applies the same meta-learning strategies to Western drum audio.
III Proposed Method
This section describes our meta-learning-based system for low-resource Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT). Fig.1 illustrates the system overview, highlighting MAML’s role in enabling few-shot adaptability. The system consists of three components: A) A CRNN-based TST model trained with MAML that transcribes audio into stroke sequences. B) A downstream $t\bar{a}la$ identification module that infers rhythmic structure from these sequences. C) Validation of the framework on a Western drums dataset for the ADT task, demonstrating adaptability across diverse drum data. We introduce two novel scoring methods for $t\bar{a}la$ identification: Needleman–Wunsch (NW) Matching Score and Stroke Ratio Score. Each component is detailed in the following subsections.
III-A Tabla Stroke Transcription (TST)
@@
This section presents our solution for TST with limited labelled data. We train and test deep learning models using three methods to identify the optimal transcription approach. Subsequently, we leverage the pre-trained model weights from the best method for MAML for TST.
III-A1 Step1: Pre-training
@@
We adopt the method of Rohit et al. [6, 7], expanding stroke classes from four to ten, with a separate CNN model for each class. Let $(X_{i},Y_{i})$ be a sample from a labelled dataset of $N$ strokes. The input $X_{i}\in\mathbb{R}^{3\times F\times T}$ is a 3-channel chunk of 15 time frames $(T)$ from log-Mel spectrograms computed using window sizes of 23ms, 46ms, and 93ms (10ms hop), stacked to combine multiple time-frequency resolutions for capturing both transient and tonal features. The model $f_{\theta}$ predicts the onset at the center frame by leveraging contextual information from surrounding frames, with $Y_{i}\in{\{0,1\}}$ indicating the presence or absence of a stroke onset.
@@
This work proposes a new mridangam and tabla stroke transcription method inspired by Sound Event Detection (SED) tasks [38]. The proposed method utilizes a CRNN for stroke transcription. The labeled dataset of $N$ samples is denoted as $D(X_{i},Y_{i})$, where $i=0,1,…,N-1$. Here, $X_{i}\in\mathbb{R}^{F\times T}$ is the input feature matrix (standardized log Mel-spectrogram), containing $F$ acoustic features extracted from each frame in each audio recording with $T$ time frames. The labels are one-hot encoded as the target output matrix $Y_{i}\in\left\{0,1\right\}^{C\times T}$, where $C$ represents the number of predefined stroke classes.
@@
After obtaining the frame-wise class predictions from the model, a post-processing step is applied to refine the output and extract onset times. This involves two stages. First, label smoothing is performed to correct short-term inconsistencies: for any frame $t\in[1,T-2]$, if the predicted class at $t$ differs from those at $t-1$ and $t+1$, but the labels at $t-1$ and $t+1$ are identical, then the label at $t$ is updated to match the surrounding frames, i.e., $Y_{c}[t]\leftarrow Y_{c}[t-1]$. Here, $T$ denotes the total number of time frames. Second, onset frames are identified as the time points where the predicted class changes between consecutive frames, i.e., $Y_{c}[t]\neq Y_{c}[t-1]$. These frame indices are then converted to onset times using the known frame rate, providing precise temporal localization of stroke events.
@@
This method is similar to the method discussed in Section III-A1, utilizing the same CRNN model architecture and training procedure. However, it adopts a different target vector methodology, similar to that in [15, 16]. The targets are represented as one-hot encoded vectors, with $Y_{i}\in\ {\{0,1\}}^{C\times T}$, where $C$ denotes the number of predefined stroke classes. In this method, only the onset frame of a particular class is assigned for that class, while all non-onset frames are assigned to the ‘No-stroke’ class. During the training phase, the model parameters are randomly initialized and updated using the stochastic gradient descent algorithm, as described by Equation (1). A weighted cross-entropy loss function is employed as indicated by Equation (2).
III-A2 Step 2: Model Agnostic Meta-learning (MAML)